DYNAMICAL FINGERPRINTING IN FORMALDEHYDE DISSOCIATION

\underline{STEVEN J. ROWLING}, SCOTT H. KABLE, School of Chemistry, The University of Sydney, Sydney, NSW, Australia, 2006; SRIDHAR A. LAHANKAR, ARTHUR G. SUITS, Department of Chemistry, Wayne State University, Detroit, MI 48202.

Photolysis of formaldehyde to H + HCO, may occur on both the electronic ground state (S_{0}) and the first excited triplet state (T_{1}). A dynamical signature distinguishing the products of these two chemical channels has been previously established.^{a} This criterion is now used to qualitatively investigate the branching ratios for these pathways at excitation energies immediately below the T_{1} potential energy barrier, and exceeding it. Comparisons are also made to the yield of the molecular product, CO, which is only generated on the S_{0} surface. Near the threshold for T_{1} participation, continual activity of both the S_{0} and T_{1} channels has been seen, with dominance fluctuating depending on the formaldehyde rovibrational state prepared.^{b} This fluctuation is confirmed in the region of the T_{1} threshold, with the triplet channel dominating the majority of states. The fluctuations diminish as the energy of the system increases away from the barrier of the triplet surface. It is also demonstrated that both pathways significantly populate the first vibrationally excited state of the HCO product once the energetic threshold is reached. Reaction on the singlet surface, however, is seen to gives rise to slightly higher vibrational excitation.

aH. M. Yin, S. H. Kable, X. Zhang and J. M. Bowman Science \underline{311}, 1443, 2006.

bL. R. Valachovic et al J. Chem. Phys. \underline{112}(6), 2752, Feb 2000.